# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

### v1

@@ 4.1 Experimental Setup
We evaluate CRAFT on three multi-hop QA benchmarks: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions Press et al. (2023); Trivedi et al. (2020). For training, we construct a mixed dataset by sampling from all three benchmarks, comprising 20,000 entries in total. For evaluation, we test on 2,000 samples from each benchmark.
@@
We compare against two model categories: (1) Open-source pre-trained models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) evaluated via in-context prompting; (2) Closed-source API models: GPT OpenAI (2025), DeepSeek DeepSeek-AI (2025), and Gemini Google DeepMind (2025). For our method, we train models at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO Shao et al. (2024).
@@
We report three evaluation metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes the token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (as detailed in Algorithm 1), providing a complementary measure beyond answer correctness.
@@
All experiments are conducted on 4 NVIDIA H100 GPUs. We use the Qwen2.5 series as base models and train with GRPO using decomposed rewards. Training hyperparameters and configurations are detailed in Appendix B.
@@ 4.2 Main Results
Table 1 presents results evaluating CRAFT against open-source pre-trained models, closed-source API models, and SFT baselines across three multi-hop QA benchmarks. Additional experiments in Appendix C.
CRAFT improves accuracy across models at all scales.
For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%. The transformation is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model—a 16 improvement, showing how CRAFT’s decomposed rewards unlock latent reasoning capacity in small models.
CRAFT improves faithfulness across models at all scales.
For the 7B model, CRAFT achieves +22.0%/+8.9%/+16.1% EM improvements over SFT on MuSiQue/HotpotQA/2WikiMHQA. While SFT often degrades Faithfulness compared to base models (50.85% vs. 58.25% for 7B on MuSiQue, a drop of -7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT. This divergence illustrates that SFT learns to imitate surface patterns, whereas CRAFT’s judge-based faithfulness reward encourages internally consistent, evidence-grounded traces.
CRAFT matches or exceeds closed-source API models.
54.35% EM on MuSiQue (vs. GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs. 61.30%/80.09%), and 78.72% EM on 2WikiMHQA (vs. 70.55%, +8.2%). This demonstrates that GRPO can train a 7B open-source model to match or exceed proprietary systems.
@@ 4.3 Ablation Study
We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation. In CRAFT, the trace template determines deterministic reward availability (, , ), so variants ablate both trace components and their corresponding rewards. Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>, disabling all deterministic rewards and judge reward. We further ablate the judge reward by training with and without . Table 2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
@@
For base models, answer-only CRAFT achieves the lowest EM (31.55%/53.40%/50.60% on MuSiQue/HotpotQA/2WikiMHQA), underperforming CRAFT (38.35%/59.30%/60.80%) by 5.9%–10.2%. While GRPO improves CRAFT substantially to 56.25%/63.25%/67.15%, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%). Notably, CRAFT cannot be evaluated for Faithfulness as it lacks reasoning traces.
Trace components reveal accuracy-faithfulness trade-offs.
For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without explicit training. After GRPO training, this gap narrows substantially (CRAFT: 80.35%/95.00%/94.80% vs. CRAFT: 86.70%/96.00%/94.10%), while CRAFTv1/CRAFTv2 improve auditability via explicit planning and citation.
CRAFT outperforms SFT across all configurations.
CRAFT achieves gains across all variants: +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models. In contrast, SFT shows frequent degradation (up to -6.6% EM on CRAFT) and systematic Faithfulness drops (-7.4% to -10.0% on MuSiQue). For CRAFT, GRPO improves MuSiQue EM by +17.5% (36.85%54.35%) while SFT decreases it by -4.5% (36.85%32.35%).
@@
Without judge reward, Faithfulness drops -8.9% to -9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable. This validates our dual reward design: deterministic rewards ensure structural correctness and answer accuracy, while judge rewards are essential for semantic faithfulness.
4.4 Training and Traces Analysis
| Question: Which member of the research project conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005 was Kevin R. Henke most critical of? |
| Gold Answer: Russell Humphreys |
| plan |
| Subquestion1: Which research project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005? |
| Subquestion2: Who was Kevin R. Henke most critical of in [ANSWER_0]? |
| gold_docs |
| [2, 7] |
| reason |
| Subquestion1: From document [2], it is clear that the RATE project was conducted by the Creation Research Society and the Institute for Creation Research between 1997 and 2005. |
| Subquestion2: Document [7] states that Kevin R. Henke has been critical of the RATE project’s results, specifically mentioning that he has argued that Russell Humphreys, a young-earth creationist involved in the project, has made errors in his research. |
| answer |
| Russell Humphreys ✓ |
4.4.1 Training Dynamics
Training performance scales with model size. Figure 3 shows total reward rising monotonically with scale: 0.5B (), 1.5B (), 3B (), and 7B (). Across all scales, reward curves converge smoothly without oscillation, indicating stable GRPO optimization under our multi-reward design. Notably, the largest relative gain appears at 1.5B, while marginal improvements diminish at 7B, suggesting saturation once most reward components approach their upper bounds.
Sub-rewards reveal fast format learning and a faithfulness threshold. Format Reward () converges within 50 training steps across all scales (0.5B: ; 7B remains near throughout), confirming that structural constraints are quickly internalized once the output template is learned. Answer Reward () scales with model capacity, reaching final scores of 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B, while Gold_doc Reward () improves steadily (0.050.27; 0.120.66; 0.190.55; 0.760.92), reflecting stronger evidence grounding in larger models.
@@ 4.4.1 Training Dynamics
Faithfulness exhibits a clear capacity threshold. Faithfulness Reward () remains near zero for 0.5B () and reaches only modest levels at 1.5B (0.10), whereas 3B and 7B show sharp improvements (, ). This transition indicates that structured consistency auditing is substantially more demanding than surface format compliance or answer correctness, and only emerges once the model has sufficient capacity to internalize and satisfy judge-based semantic constraints.
4.4.2 Output Traces Case Study
We analyze the quality of generated reasoning traces to understand how CRAFT structures multi-hop reasoning. Table 3 presents a representative example using the CRAFT, demonstrating CRAFT’s systematic approach with explicit planning, gold document citation, step-by-step reasoning, and final answer extraction. Additional case study in Appendix A. About traces format analysis in Appendix C.3.
@@ 5 Conclusion
We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA. CRAFT employs a dual reward design that combines deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness. Experiments on three benchmarks show that CRAFT consistently improves both accuracy and faithfulness over base models and SFT, with CRAFT matching or exceeding closed-source API models. Ablation results highlight the importance of structured traces and judge rewards for learning faithful reasoning, while training dynamics analysis indicates that faithfulness learning requires sufficient model capacity.
Appendix A Faithfulness Case Study
We present two representative cases from MuSiQue comparing Qwen2.5-7B Base and CRAFT outputs, demonstrating how CRAFT training improves both answer correctness and reasoning faithfulness. Each case includes the complete judge evaluation across all four faithfulness dimensions defined in Algorithm 1.
A.1 Case 1: Base (EM=0, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates how an unfaithful and incorrect response is transformed into a faithful and correct one.
Gold Answer: Mario Andretti
| Aspect | Base Model (EM=0, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Where is the Oh Yeah album from?
Subq1: Where is the Oh Yeah performer from?
Subq2: What is the largest city and capital of the state where the Oh Yeah performer is from? Subq3: Who won the Indy car race in the largest city and capital of that state? |
| Gold Docs | [5, 12, 14] | [11, 12, 1] |
| Reason | [Answer1] Charles Mingus, the creator of the Oh Yeah album, was born in Nogales, Arizona (doc [5]). [Answer2] The largest city and capital of Arizona is Phoenix (doc [12]). [Answer3] The Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and the winner of the Indy car race there in 2016 was Ryan Hunter-Reay (doc [1]). | Doc [11] indicates that the Oh Yeah performer, Charles Mingus, is from Nogales, Arizona. Doc [12] states that the capital and largest city of Arizona is Phoenix. Doc [1] mentions that the Desert Diamond West Valley Phoenix Grand Prix was held in Phoenix, and it is an Indy car race where Mario Andretti won. |
| Answer | Ryan Hunter-Reay | Mario Andretti ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 1 - Plan steps followed in order | 1 - Plan followed: performer→state→city→winner |
| Receipt-Reason () | 0 - Cites doc [1] not in gold_docs [5,12,14] | 1 - All cited docs [11,12,1] in gold_docs |
| Reason-Answer () | 1 - Answer matches reasoning conclusion | 1 - “Mario Andretti” supported by reasoning |
| Evidence Grounded () | 0 - “Ryan Hunter-Reay won” not in doc [1] | 1 - All claims verified in cited docs |
| Overall () | 0.0 - Failures: doc_outside_gold_docs, evidence_not_supported | 1.0 - All checks passed |
A.2 Case 2: Base (EM=1, Faith=0) CRAFT (EM=1, Faith=1)
This case demonstrates that correct answers can arise from unfaithful reasoning, and CRAFT enforces faithfulness even when accuracy is preserved.
Question: What is the average yearly income for a person of the same nationality as the creator of MacGruder and Loud?
Gold Answer: $72,641
| Aspect | Base Model (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
| Plan |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the nationality of [Answer1]? Subq3: What is the average yearly income for a person of [Answer2] nationality? |
Subq1: Who is the creator of MacGruder and Loud?
Subq2: What is the average yearly income for a person of the same nationality as the creator? |
| Gold Docs | [4, 5, 2] | [8, 4] |
| Reason | [Answer1]: From document 4, we learn that the creator of MacGruder and Loud is Aaron Spelling. Document 5 tells us that Lorraine Crosby, known as “Mrs. Loud,” is English. Therefore, the nationality of the creator of MacGruder and Loud is English. Document 2 provides the average yearly income for Americans, which is the closest nationality we have data for. The mean household income in the United States was $72,641. | Doc [8] states that MacGruder and Loud was one of the few failures from Aaron Spelling’s production company. Doc [4] provides information about Aaron Spelling, indicating that he was an American film and television producer. Therefore, the creator of MacGruder and Loud is American. The average yearly income for an American, as mentioned in doc [2], is $72,641. |
| Answer | $72,641 ✓ (correct answer, but unfaithful reasoning) | $72,641 ✓ |
| Judge Evaluation | ||
| Plan-Reason () | 0 - Identifies Aaron Spelling but concludes “English” | 1 - Plan followed: creator→nationality→income |
| Receipt-Reason () | 0 - Misuses doc [5] (Lorraine Crosby creator) | 1 - All cited docs [8,4] in gold_docs |
| Reason-Answer () | 0 - Derives “English” but uses US income | 1 - “$72,641” supported by reasoning |
| Evidence Grounded () | 0 - “Creator is English” unsupported | 1 - Spelling→American→US income verified |
| Overall () | 0.0 - Failures: plan_ignored, answer_not_supported, evidence_not_supported | 1.0 - All checks passed |
A.3 Summary
These cases illustrate two critical aspects of CRAFT’s faithfulness enforcement:
Case 1 shows that CRAFT corrects both unfaithful reasoning (citing documents outside gold_docs, hallucinating unsupported claims) and incorrect answers.
Case 2 demonstrates that answer correctness does not imply reasoning faithfulness. The Base model produces the correct answer through a flawed reasoning chain (confusing Lorraine Crosby with Aaron Spelling), while CRAFT ensures the answer is derived through a consistent, evidence-grounded process.
The judge evaluation tables reveal that CRAFT training optimizes all four faithfulness dimensions simultaneously: plan-reason consistency (), receipt-reason consistency (), reason-answer consistency (), and evidence grounding (). This multi-dimensional supervision prevents shortcut learning and ensures that models produce not just correct answers, but trustworthy reasoning traces.
Appendix B Implementation Details
@@
We implement CRAFT using MS-Swift Zhao et al. [2025], an efficient and scalable fine-tuning framework for large language models. MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning. All experiments are conducted on NVIDIA H100 96GB x4 GPUs.
@@
We train four model scales based on the Qwen2.5-Instruct series: 0.5B, 1.5B, 3B, and 7B parameters. All models use bfloat16 precision with Flash Attention 2 for memory-efficient training. Table 6 summarizes the key hyperparameters.
@@
For the faithfulness reward and evaluation, we employ Qwen3-Max Team [2025] as the LLM judge. This model provides reliable semantic consistency assessment while maintaining computational efficiency. Because Qwen3-Max already functions as the judge for reward computation and evaluation, it is excluded from all performance comparison experiments to avoid role overlap.
@@
The reward functions vary across prompt templates based on their structural components:
CRAFT/CRAFT: , , (includes <gold_docs> field)
CRAFT/CRAFT: , (no explicit document citation)
CRAFT: , (direct answer generation)
All reward weights are set to 1.0, treating each component equally. The faithfulness reward is computed via LLM judge during evaluation but not used as a training signal in the ablation experiments (Table 2, “w/o judge reward”).
@@
Each prompt template uses a dedicated training dataset of 20,000 samples, formatted according to the template’s structural requirements. The data distribution is: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA samples. Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns that indicate diverse linguistic characteristics across the three multi-hop QA datasets.
Appendix C Additional Experiments
This section presents comprehensive results for prompt template variants CRAFT–CRAFT, including training dynamics and evaluation results across all model scales.
C.1 Training Dynamics
Figures 5–8 illustrate the training dynamics for CRAFT–CRAFT templates. Consistent with CRAFT (main paper Figure 3), all templates exhibit smooth convergence with larger models achieving higher total rewards.
Template-specific patterns: (1) CRAFT (Figure 5) shows similar dynamics to CRAFT, with all four reward components (, , , ) contributing to optimization. (2) CRAFT/CRAFT (Figures 6–7) lack but maintain stable training through format and answer rewards. (3) CRAFT (Figure 8) uses only and —notably, the 0.5B model shows minimal improvement, foreshadowing its evaluation failure.
Faithfulness reward verification.
To verify that the low training faithfulness reward (Figure 3) reflects optimization dynamics rather than learning failure, we measured the four faithfulness metrics (Algorithm 1) on test samples of the Qwen2.5-1.5B CRAFT model under greedy decoding (temp=0.0). Among test samples with correct answers (EM=1), the four faithfulness metrics achieve 76.85%, 68.95%, 70.15%, and 61.35% respectively, with an arithmetic mean of 69.33%. This indicates that while 69.7% of correct predictions exhibit fully faithful reasoning, 30.3% still fail at least one faithfulness check, demonstrating that answer correctness alone does not guarantee reasoning faithfulness. This test-time value is approximately six times higher than the training-time average of 0.11 observed for this model (Figure 3), confirming that GRPO’s on-policy sampling at high temperature (temp=1.0) generates diverse outputs including many low-quality traces during exploration, resulting in lower average rewards during training, while the learned policy concentrates probability mass on high-quality outputs under greedy decoding at test time.
@@ C.1 Training Dynamics
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team [2024] | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
| Qwen2.5-1.5B Team [2024] | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
| Qwen2.5-3B Team [2024] | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
| Qwen2.5-7B Team [2024] | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
| Qwen3-4B Team [2025] | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
| Qwen3-30B Team [2025] | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
| API | ||||||||||
| GPT-5-mini OpenAI [2025] | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
| Deepseek-V3.2 DeepSeek-AI [2025] | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
| Gemini-2.5-Flash Google DeepMind [2025] | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team [2024] | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
| Qwen2.5-1.5B Team [2024] | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
| Qwen2.5-3B Team [2024] | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
| Qwen2.5-7B Team [2024] | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
| Ours (RL) | ||||||||||

### v2

@@ 4.1 Experimental Setup
We evaluate CRAFT on three multi-hop QA benchmarks: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). These datasets represent diverse multi-hop reasoning types including bridge reasoning, comparison, and compositional questions. For training, we construct a mixed dataset of 20,000 entries sampled from all three benchmarks. For evaluation, we test on 2,000 samples from each benchmark.
@@
We compare against: (1) Open-source models: Qwen2.5 series (0.5B, 1.5B, 3B, 7B) Team (2024) and Qwen3 series (4B, 30B) Team (2025) via in-context prompting; (2) Closed-source API models: GPT-5-mini OpenAI (2025), DeepSeek-V3.2 DeepSeek-AI (2025), and Gemini-2.5-Flash Google DeepMind (2025). We train CRAFT at four scales (0.5B, 1.5B, 3B, 7B) using both SFT and GRPO.
@@
We report three metrics: (1) Exact Match (EM) measures whether the predicted answer exactly matches the gold answer after normalization; (2) F1 score computes token-level overlap between prediction and gold answer; (3) Faithfulness evaluates reasoning quality using an LLM judge that audits internal consistency and evidence grounding (Algorithm 1).
@@ 4.2 Main Results
Table 3.3.2 presents results across three multi-hop QA benchmarks.
CRAFT improves accuracy across all model scales.
For the 7B model, CRAFT achieves gains of +17.5%/+9.7%/+21.0% EM on MuSiQue/HotpotQA/2WikiMHQA compared to base Qwen2.5-7B, corresponding to relative improvements of 47%/17%/36%. The improvement is more dramatic at smaller scales: CRAFT achieves 24.80% EM on HotpotQA compared to 1.55% for the base model, a 16 improvement, demonstrating that CRAFT’s decomposed rewards are effective even at very small scales.
CRAFT improves faithfulness across all model scales.
While SFT often degrades Faithfulness compared to base models (e.g., 50.85% vs. 58.25% for 7B on MuSiQue, a drop of 7.4%), CRAFT improves accuracy and faithfulness simultaneously, achieving +29.5%/+11.5%/+15.9% in Faithfulness over SFT.
CRAFT matches or exceeds larger closed-source API models.
CRAFT achieves 54.35% EM on MuSiQue (vs. GPT-5-mini 53.45%), 66.61%/82.09% EM/F1 on HotpotQA (vs. 61.30%/80.09%), and 78.72% EM on 2WikiMHQA (vs. 70.55%, +8.2%). This demonstrates that GRPO with decomposed rewards can train a 7B open-source model to match or exceed proprietary systems.
@@ 4.3 Ablation Study
We conduct two ablation studies: (1) trace template ablation with aligned deterministic rewards, and (2) judge reward ablation. Starting from the full template CRAFT (<plan>+<gold_docs>+<reason>+<answer>), we progressively remove components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains <reason>+<answer>, and CRAFT retains only <answer>. Table 3.3.2 reports results for Qwen2.5-7B across Base, SFT, CRAFT without judge reward, and full CRAFT.
@@
Answer-only CRAFT achieves the lowest base EM (31.55%/53.40%/50.60%), underperforming CRAFT by 5.9%–10.2%. While GRPO improves CRAFT substantially, structured variants still achieve higher performance (e.g., CRAFT: 58.35%/67.28%/79.70%).
Simpler traces achieve higher base faithfulness but lower auditability.
For base models, simpler CRAFT yields higher Faithfulness (76.60%/93.75%/86.15%) than full CRAFT (58.25%/83.65%/76.85%), as additional fields introduce inconsistency opportunities without training. After GRPO, this gap narrows substantially, while CRAFT/CRAFT improve auditability via explicit planning and citation.
SFT degrades faithfulness despite improving format compliance.
CRAFT achieves +7.2% to +24.7% EM and +2.3% to +22.1% Faithfulness over base models, while SFT shows frequent degradation (up to 6.6% EM, 10.0% Faithfulness on MuSiQue).
@@
Without judge reward, Faithfulness drops 8.9% to 9.9% on MuSiQue across CRAFT–CRAFT, while EM scores remain comparable. This validates our dual reward design: deterministic rewards ensure structural correctness, while judge rewards are essential for semantic faithfulness. Notably, adding the judge reward also improves EM (e.g., +3.0% on CRAFT MuSiQue), confirming that faithful reasoning supervision enhances rather than compromises answer accuracy.
4.4 Training Dynamics
Figure 3 shows the training dynamics of GRPO across model scales using CRAFT(full format).
Training performance scales with model size.
Total reward increases monotonically with scale, and all curves converge smoothly, indicating stable optimization under our multi-reward design. Larger models achieve both higher starting points and higher ceilings, with the 7B model reaching a total reward of 3.41 compared to 1.47 for 0.5B.
Format compliance is learned quickly; faithfulness requires sufficient model capacity.
@@ 4.4 Training Dynamics
converges within 50 steps across all scales, indicating that structural constraints are the easiest to satisfy. scales with capacity, reaching 0.22/0.62/0.63/0.89 for 0.5B/1.5B/3B/7B; notably, 1.5B and 3B achieve similar answer accuracy but diverge sharply on faithfulness. remains near zero for 0.5B and reaches only modest levels at 1.5B, whereas 3B and 7B show clear improvements (, ). This indicates that faithfulness learning is substantially more demanding than format compliance or answer accuracy, and only emerges with sufficient model capacity.
CRAFT achieves near-perfect format compliance across all scales.
Full format compliance results across all variants are reported in Appendix B.4.
| Aspect | Both | Both | LLM | LLM | Agr. | |
| Pass | Fail | Human | Human | |||
| Plan-Reason | 393 | 78 | 15 | 14 | 94.2% | 0.81 |
| Evidence Gr. | 375 | 83 | 25 | 17 | 91.6% | 0.75 |
| Answer Deriv. | 410 | 71 | 9 | 10 | 96.2% | 0.86 |
| Gold Doc Cit. | 370 | 80 | 33 | 17 | 90.0% | 0.70 |
| Overall (Avg.) | 387 | 78 | 20 | 15 | 93.0% | 0.78 |
4.5 Human Validation of Judge Reliability
To validate our LLM-based faithfulness metric, two annotators with NLP research experience independently assessed 500 randomly sampled MuSiQue reasoning traces from CRAFT using the same four-aspect binary criteria as the LLM judge (Algorithm 1).
We select MuSiQue as it is the most challenging benchmark, requiring up to four reasoning hops. Results show 93.0% overall agreement with Cohen’s Artstein (2017), indicating substantial agreement (Table 3). Answer Derivation achieves the highest agreement (), while Gold Document Citation shows lower agreement (), reflecting inherent ambiguity in whether cited documents provide essential versus supplementary support. These results confirm that the LLM judge aligns well with human judgment in absolute terms. We further re-evaluate all configurations with an independent judge (Gemini-2.5-Flash) in Appendix C; model rankings are perfectly preserved (Spearman = 1.0). Together, these two evaluations address the circularity of using Qwen3-Max for both training and evaluation: human validation confirms that the judge’s criteria reflect genuine faithfulness rather than idiosyncratic preferences, while cross-judge consistency confirms that the reported improvements are not artifacts of a specific evaluator.
| Aspect | Base (EM=1, Faith=0) | CRAFT (EM=1, Faith=1) |
[…]
Subq1: Who is the creator?
Subq2: What is the nationality? Subq3: What is the avg income? |
Subq1: Who is the creator?
Subq2: What is the avg income for same nationality? |
[…]
| 0 - Concludes “English” despite identifying Spelling | 1 - creatornationalityincome | |
| 0 - Misuses doc [5] (Crosby creator) | 1 - All cited docs in gold_docs | |
| 0 - Derives “English” but uses US income | 1 - $72,641 supported by reasoning | |
| 0 - “Creator is English” unsupported | 1 - All claims verified | |
| 0.0 | 1.0 | |
4.6 Case Study
Table 4 presents a case from MuSiQue where the base model produces the correct answer through unfaithful reasoning (EM=1, =0), while CRAFT achieves both correctness and faithfulness (EM=1, =1). The base model answers “$72,641” correctly but confuses Aaron Spelling with Lorraine Crosby from a distractor document, concluding the creator is “English” yet using US income data. This is a typical distractor-induced entity confusion: the model latches onto a surface-similar passage and constructs a plausible-sounding but internally inconsistent chain. Notably, all four judge checks fail (====0), demonstrating how the chain of faithfulness catches errors that outcome-level metrics miss entirely. In contrast, CRAFT correctly traces SpellingAmericanUS income, citing only relevant documents and passing all four faithfulness checks. An additional case study where both accuracy and faithfulness are corrected is provided in Appendix D.
@@ 5 Conclusion
We present CRAFT, a GRPO-based reinforcement learning framework that trains RAG-based LLMs to generate structured, machine-auditable reasoning traces for multi-hop QA. By combining deterministic rewards for structural correctness with judge-based rewards for semantic faithfulness, CRAFT consistently improves both accuracy and faithfulness across model scales, with CRAFT matching or exceeding closed-source API models. Notably, semantic judge-based rewards improve answer accuracy rather than compromise it.
Limitations
Our human validation of judge reliability is conducted on 500 samples from MuSiQue only. While we select MuSiQue as the most challenging benchmark and supplement with cross-judge evaluation (Appendix C), extending human annotation to all three datasets would further strengthen the reliability analysis; we leave this to future work due to the cost of expert annotation on multi-hop reasoning traces. Second, due to the substantial computational cost of GRPO training with online judge inference, our experiments are limited to models up to 7B parameters trained on 20K samples from three multi-hop QA benchmarks (HotpotQA, 2WikiMHQA, MuSiQue). A central finding of this work is that faithfulness supervision can improve rather than compromise answer accuracy. We focus on establishing this finding within the training distribution first, as confirming that the accuracy–faithfulness synergy holds under controlled conditions is a necessary prerequisite before testing its generalization. Scaling to larger models and evaluating whether this synergy transfers to out-of-distribution QA domains remains important future work.
Appendix A Implementation Details
This section expands on the experimental setup described in §4.
@@ Appendix A Implementation Details
We implement CRAFT using MS-Swift Zhao et al. (2025), an efficient fine-tuning framework for large language models. MS-Swift provides native support for GRPO training with vLLM acceleration, enabling efficient on-policy sampling during reinforcement learning. All experiments are conducted on NVIDIA H100 96GB 4 GPUs. Table A summarizes the key hyperparameters.
@@
We train four model scales based on the Qwen2.5-Instruct series: 0.5B, 1.5B, 3B, and 7B parameters. All models use bfloat16 precision with Flash Attention 2 for memory-efficient training. For SFT baselines, we fine-tune each model on the same training data using standard cross-entropy loss with a learning rate of 2e-5 and 3 epochs.
@@
We employ Qwen3-Max Team (2025) as the LLM judge for both the faithfulness reward during training and the Faithfulness metric during evaluation. The judge is prompted with the query, retrieved documents, and the model’s trace, and returns binary scores for each applicable audit dimension. We use temperature 0.0 for deterministic judge outputs. Because Qwen3-Max serves dual roles as reward provider and evaluator, it is excluded from all performance comparisons. We verify that this dual usage does not bias our conclusions through cross-judge evaluation with Gemini-2.5-Flash (Appendix C).
@@
The active reward components vary across trace templates based on the available structural fields: CRAFT/CRAFT: , , , (full supervision); CRAFT/CRAFT: , , (no citation reward as <gold_docs> is absent); CRAFT: , only (no reasoning fields for faithfulness audit). All reward weights are set to 1.0. This is a natural default since , , , and all produce values in ( is binary, and are F1 scores, and is an average of binary checks), so equal weighting does not introduce scale imbalance. Moreover, converges to 1.0 within 50 training steps across all scales (Figure 3), contributing negligible gradient signal thereafter. In the ablation without judge reward (Table 3.3.2, “w/o judge reward”), is removed from training but still computed during evaluation.
@@
Each prompt template uses a dedicated training dataset of 20,000 samples: 10,000 MuSiQue, 5,000 HotpotQA, and 5,000 2WikiMHQA. We oversample MuSiQue as it contains questions requiring 2–4 reasoning hops, providing greater diversity in reasoning depth compared to the predominantly 2-hop questions in HotpotQA and 2WikiMHQA. Figure 4 visualizes the training data distribution using t-SNE, revealing distinct clustering patterns across the three datasets.
@@
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.05 | 1.06 | 1.35 | 1.15 | 3.91 | 4.20 | 1.90 | 5.13 | 2.25 |
| Qwen2.5-1.5B | Pre-Trained | 3.95 | 7.70 | 7.20 | 16.95 | 25.12 | 21.45 | 15.50 | 21.53 | 18.15 |
| Qwen2.5-3B | Pre-Trained | 21.75 | 31.54 | 42.25 | 49.65 | 63.82 | 74.75 | 43.35 | 52.51 | 57.95 |
| Qwen2.5-7B | Pre-Trained | 34.35 | 44.62 | 65.45 | 54.80 | 69.75 | 88.05 | 56.05 | 65.68 | 77.20 |
| Qwen3-4B | Pre-Trained | 34.35 | 43.30 | 57.45 | 59.65 | 75.25 | 87.95 | 63.20 | 72.63 | 82.05 |
| Qwen3-30B | Pre-Trained | 48.15 | 57.82 | 71.80 | 63.80 | 78.38 | 89.60 | 69.95 | 78.35 | 88.25 |
[…]
| GPT-5-mini | API | 51.70 | 67.02 | 88.30 | 61.45 | 80.10 | 96.90 | 69.95 | 80.64 | 94.00 |
| Deepseek-V3.2 | API | 52.05 | 67.62 | 85.40 | 62.20 | 79.42 | 96.40 | 69.65 | 79.02 | 97.15 |
| Gemini-2.5-Flash | API | 50.75 | 61.95 | 81.40 | 63.35 | 79.09 | 94.80 | 70.20 | 79.31 | 94.20 |
[…]
| Qwen2.5-0.5B | Trained | 0.00 | 0.55 | 0.70 | 0.60 | 2.10 | 2.35 | 1.05 | 2.75 | 1.20 |
| Qwen2.5-1.5B | Trained | 4.50 | 7.64 | 7.80 | 17.85 | 26.71 | 36.20 | 27.70 | 34.14 | 38.05 |
| Qwen2.5-3B | Trained | 22.25 | 31.55 | 36.95 | 50.60 | 64.54 | 78.55 | 52.05 | 61.39 | 72.25 |
| Qwen2.5-7B | Trained | 29.85 | 41.17 | 56.10 | 55.70 | 70.25 | 86.70 | 59.25 | 67.73 | 78.30 |
[…]